OpenStudy (australopithecus):

computing the limit of the series \[\lim_{n \rightarrow \infty} \frac{n^{3} + \sin(-2011n)}{n^{3} + 110111}\] = DNE is this correct?

OpenStudy (australopithecus):

My reasoning is \[\lim_{n \rightarrow \infty} \frac{\sin(-2011n)}{n^{3}+110111} = DNE\]

OpenStudy (australopithecus):

Thus the limit cannot be computed because sin(infinity) = 1,0

OpenStudy (across):

Hint:\[|\sin x|\leqslant1\text{, }\forall x\in\mathbb{R}\]

OpenStudy (valpey):

@across makes a good point. This looks like a job for L'Hospital.

OpenStudy (across):

It's actually pretty straightforward.

OpenStudy (australopithecus):

I thought you couldn't compute the limit as x goes to infinity of trig functions

OpenStudy (australopithecus):

well at least for sin(x) as it goes to 1 and 0

OpenStudy (australopithecus):

thus a limit is impossible to compute

OpenStudy (across):

Never mind: you edited the OP.

OpenStudy (australopithecus):

yeah well it would be 1 + DNE = DNE

OpenStudy (across):

In your new case, break the fraction into two parts and evaluate each individually.

OpenStudy (australopithecus):

I did

OpenStudy (across):

I still don't understand your reasoning behind the DNE.

OpenStudy (ash2326):

We don't need to use L'hospital :) \[\lim_{n \rightarrow \infty} \frac{n^{3} + \sin(-2011n)}{n^{3} + 110111}\] divide numerator and denominator by n^3 \[\huge\lim_{n \rightarrow \infty} \frac{1 + \frac{\sin(-2011n)}{n^3}}{1 + \frac{110111} {n^3}}\] we know \[|\sin x|\le 1\] so a finite quantity by an infinite quantity is zero so \[\huge\lim_{n \rightarrow \infty}\frac{1+0}{1+0}\] hence the limits exist

OpenStudy (across):

@ash2326 illustrated my point exactly.

OpenStudy (australopithecus):

My reasoning behind it is that sin(x) has two possible numbers it can be 1 and 0 thus the limit is not determinate.

OpenStudy (australopithecus):

as it approaches infinity

OpenStudy (across):

No, sin can be anything in the real domain \([-1,1]\).

OpenStudy (australopithecus):

oh sorry you are correct, brain fart http://www.wolframalpha.com/input/?i=lim+n-%3Einfinity+sin%28n%29

OpenStudy (across):

What you should understand here is that you're basically dividing a number between \(-1\) and \(1\) by infinity.

OpenStudy (australopithecus):

still how can you determine a limit if there is no one number

OpenStudy (ash2326):

@Australopithecus we know that it's between -1, 1 so we know it's a finite quantity. Hence \[\lim_{n\to \infty} \frac{\sin n}{n}=0\] if we had \[\lim_{n\to \infty} {\sin n}\] then it'd not exist

OpenStudy (australopithecus):

oh alright so sin(n)/n as n goes to infinity = 0

OpenStudy (australopithecus):

thanks for clarifying that for me

OpenStudy (australopithecus):

so the answer is 1

OpenStudy (australopithecus):

or 1 + 0